2+4+7+11+16+......to n terms= What us the sum total

Grade:12th pass

1 Answers

Sanjna Subramanian
6 years ago
sum of n terms = n/2 [2a+d(n-1)]                                                                                                                                                                                                          =n/2[2*2+2(n-1)]                                                                                                                                                                                                             =n/2(2+2n)                                                                                                                                                                                                                  =n(sq.) +n                                                                                                                                                                                                therefore sum of n terms of the sequence is n(sq.) +n